ingrid wants to buy car in 7 years. How much money must she deposit at the end of each quarter in an account paying 5.1% compunded quarterly so that she will have enough to pay for her car?

wants to buy a $16,000 car
Answer #1

Future value (FV) = $16,000
Nper = 7 * 4 = 28
Rate = 5.1%/4 = 1.275%
PV = 0

Amount of deposit at the end of each quarter is calculated by using the following Excel formula:
=PMT(rate,nper,pv,fv)
=PMT(1.275%,28,0,-16000)
= $479.08 (approx).

Amount of deposit at the end of each quarter is $479.08 (approx).